# Are Vision xLSTM Embedded $U$-Nets More Reliable in Medical 3D Image Segmentation?

### v2

@@ 3.2 Feature reconstruction path
Input: Training dataset $\Delta=\{(I_{j},O_{j})\}_{j=1}^{m}$, Patch size $P$, Embedding dimension $Z$, Number of Vision-xLSTM blocks $L$, Projection matrix $K$, Positional embedding $K_{pos}$, learning rate $\epsilon$ and number of iterations $\iota$
@@
20:     for $i=1$ to $L$ do /* Apply Vision-xLSTM Blocks */
21:         Compute $Q^{i},K^{i},V^{i}$ matrices and attention scores $S^{i}$ with a causal mask
22:         Update memory state $\tilde{h}^{i}_{t}$ and obtain output $o^{i}$
@@ 4 Experimental Details
This section describes the loss functions and performance metrics used for the evaluation, details of the datasets used, and the experimental results.
@@ 4.1 Loss function and metrics
A loss function combination of Dice $\mathcal{L}_{dice}=\gamma-\sum_{\gamma=1}^{c}\left(\frac{2\sum_{i=1}^{T}\hat{z}_{\gamma,i}z_{\gamma,i}+\mu}{\sum_{i=1}^{T}\hat{z}_{\gamma,i}+z_{\gamma,i}+\mu}\right)$ [5] and Categorical Cross Entropy [32] $\mathcal{L}_{cce}=-\frac{1}{T}\sum_{i=1}^{T}\sum_{\gamma=1}^{c}z_{\gamma,i}\log{(\hat{z}_{\gamma,i})}$, was employed. Here, $c$ represented the total number of classes to be predicted, $\hat{z}_{\gamma,i}$ and $z_{\gamma,i}$ corresponded to the predicted and ground truth values (respectively) for the $i^{th}$ voxel for class $\gamma$, $T$ denoted the total number of voxels in the input, and $\mu$ was the additive smoothing parameter used to overcome division-by-zero errors. The composite loss function $\mathcal{L}$ was expressed as
@@
Experimental Setup: The PyTorch and MONAI frameworks were used to implement and train U-VixLSTM in Python 3.9 on an NVIDIA RTX A6000 GPU with 48GB of RAM. AdamW Optimizer was used, with a learning rate of 1e-4 and a weight decay of 1e-5.
Figure 2: Distribution of (a) various abdominal organs, (b) skin lesion surface area, and (c) cardiac anatomical structures, among different patients in the Synapse, ISIC, and ACDC datasets, respectively.
@@ 4.2 Datasets
The data set Synapse comprises 30 CT volumes, with sizes ranging from 512 $\times$ 512 $\times$ 85 to 512 $\times$ 512 $\times$ 198. CT volumes are manually annotated by experts in the field to highlight the different abdominal organs. A model is trained to segment nine distinct organs of the abdominal cavity viz. spleen, left kidney, right kidney, liver, gall bladder, pancreas, stomach, right adrenal gland, and left adrenal gland. The spleen, liver, and stomach are classified as larger organs, while the kidneys, gall bladder, pancreas, and adrenal glands are smaller in size. The CT volumes were divided into training and test sets, with 80% used for training and the remaining 20% used for testing to evaluate the generalization performance of the model. The voxel intensities were restricted between -170 HU and 250 HU, followed by normalization of the intensity to the range of [0,1]. Data augmentation was effective in expanding the size of the training dataset.
@@
Fig. 2 illustrates the distribution of abdominal and cardiac organ volumes, as well as the surface area of skin lesions, between patients in the three datasets. The plots demonstrate substantial evidence of inter-patient variability across the datasets, indicating significant differences in the sizes of various organs or lesion areas among patients. Quite a significant difference in morphology is evident in the spleen, liver, and stomach, as shown in Fig. 2(a), in the cardiac organs of Fig. 2(b), and in the total surface area of the lesion illustrated in Fig. 2(c), as attributed to long tails and broader distributions. The median values indicate significant heterogeneity in the average volumes of the various abdominal organs. The liver exhibits the largest average volume, followed by the stomach and spleen. The adrenal gland exhibits the lowest average volume among the glands. The diverse sizes of target structures hinder the ability of the model to generalize across datasets. Fig. 2(c) illustrates a right-skewed distribution of the surface area of the lesions. This suggests that patients generally have smaller lesion areas compared to larger ones. This may indicate a class imbalance within the dataset.
@@
4.3 Results
Table 1: Comparison of different variants of U-VixLSTM, with increasing number of Vision-xLSTM blocks (×$\times L$) and increasing number of convolution layers, on Synapse data. Best results and the selected model configuration are highlighted in bold.
@@ 4.3 Results
Table 2: Comparison with state-of-the-art models, on multi-organ segmentation (Synapse) dataset. The best results are highlighted in bold.
@@
$U$-Net [3] 0.9112 0.9007 0.9181 0.5645 0.9572 0.6967 0.78 0.616 0.5872 0.7702 0.6667 38.83
$V$-Net [5] 0.8874 0.9251 0.9244 0.5858 0.9487 0.7511 0.7953 0.6004 0.4497 0.7631 0.6537 25.29
$U$-Net++ [4] 0.9118 0.9196 0.8905 0.6921 0.9524 0.7536 0.7869 0.5245 0.1032 0.7261 0.6254 55.27
Attention $U$-Net [6] 0.9109 0.877 0.872 0.5835 0.9585 0.5566 0.7846 0.5991 0.5398 0.7424 0.6262 51.61
$U$-Net + EfficientNet-b0 [33] 0.8541 0.8919 0.8804 0.637 0.9077 0.5018 0.7097 0.6048 0.5195 0.723 0.6056 67.47
$U$-Net + EfficientNet-b1 [33] 0.7414 0.8059 0.8256 0.55 0.9019 0.5416 0.616 0.525 0.5035 0.6679 0.5539 75.85
$U$-Net 3+ [7] 0.8736 0.9229 0.8584 0.6659 0.9383 0.6823 0.7259 0.5657 0.4764 0.7455 0.6254 54.43
SwinUNet [14] 0.9482 0.93 0.9245 0.7617 0.9622 0.8046 0.8059 0.689 0.6159 0.8269 0.7261 13.99
UNETR [13] 0.8951 0.9055 0.895 0.5274 0.9461 0.6668 0.7847 0.5268 0.4978 0.7384 0.6284 26.94
TransUNET [8] 0.852 0.8828 0.785 0.7608 0.587 0.7218 0.8773 0.5876 0.4672 0.7246 0.6679 36.81
TransAttUNET [15] 0.9045 0.8761 0.916 0.4958 0.9408 0.5997 0.6682 0.6559 0.5948 0.7391 0.658 15.82
Swin UMamba [30] 0.8778 0.9037 0.9112 0.4206 0.9161 0.6238 0.6527 0.589 0.5617 0.7174 0.6334 20.1
UNETR++ [16] 0.8061 0.8050 0.8201 0.5208 0.8868 0.5249 0.5576 0.589 0.4577 0.6631 0.534 8.61
DS-TransUNET [17] 0.4965 0.6113 0.5252 0.5397 0.679 0.5295 0.5857 0.5745 0.5266 0.5631 0.4644 14.79
U-VixLSTM 0.95 0.9371 0.9366 0.8104 0.9635 0.7878 0.8304 0.6709 0.6458 0.8318 0.7286 4.8
Table 1 presents an ablation study on Synapse data, demonstrating the impact of different numbers ($L$) of Vision-xLSTM blocks and a varying number of convolution layers along the feature extraction path. The average scores $DSC$, $HD95$, and $IoU$ are reported for the different organs. The ablation study reflects the trade-off between the complexity of the model and its performance. Increasing the number of convolution layers from three to four shows a significant gain in metric scores. However, the performance saturates on further increase in the number of layers. The best $DSC$ and $IoU$ values were observed for the lowest number of Vision-xLSTM blocks. The $HD95$ scores showed minimal improvement with 18 Vision-xLSTM blocks, compared to six. This implies that the lower computational and memory requirements of U-VixLSTM resulted in higher efficiency w.r.t. $DSC$ and $IoU$. However, a slightly deeper architecture may provide small improvements in accurately defining object boundaries with increased computational overhead. It was seen that a model with 6 Vision-xLSTM blocks attained an offset between efficient performance and low computational complexity.
@@
The performance of our U-VixLSTM was next compared with that of the state-of-the-art (SOTA) algorithms in the literature in terms of metrics $DSC$, $IoU$, $HD95$ on Synapse, ISIC and ACDC datasets. Table 2 presents a comprehensive breakdown of performance in the different organs (in Synapse data) in the context of DSC. The average results for HD95 and IoU are provided for the nine abdominal organs. U-VixLSTM is found to outperform other SOTA, with respect to the mean DSC, IoU and HD95, with scores of 83.18%, 72.86%, and 4.8, respectively. Our proposed U-VixLSTM demonstrated superior performance in generating the highest $DSC$ values for the segmentation of larger organs (such as the spleen, liver) and smaller organs (such as the kidneys, pancreas, gall bladder and left adrenal gland). Performance remained stable, despite the reduction in organ size, as indicated by the consistently high $DSC$ scores observed in smaller and larger organs. This illustrates the robustness of our model in utilizing acquired knowledge on anatomical structures with varying shapes and sizes. It is found to precisely identify and define the target areas possessing unique structures.
@@
Figure 3: Comparative performance of U-VixLSTM and other baseline architectures, on the Synapse dataset, through sample segmentation maps. The (a) input CT image, (b) corresponding ground truth, with the respective output from (c) U-VixLSTM, (d) Swin UNETR, (e) UNETR, (f) V-Net, and (g) TransAttUNet.
@@
Fig. 3 shows the sample segmentation output of U-VixLSTM and other architectures, such as Swin-UNETR, UNETR, $V$Net, and TransAttUNet, on the Synapse dataset. It is evident that the U-VixLSTM model demonstrates a significantly higher level of similarity to the ground truth, compared to the maps produced by these other baseline models. The results of the Swin UNETR and UNETR models show $FP$ regions corresponding to the stomach, in the first row of Figs. 3 (d) and (e). The result of $V$-Net in the upper row of Fig. 3 (f) demonstrates that pixels associated with the boundary areas of the spleen were incorrectly classified. It also demonstrates a lack of precision in identifying the anatomical regions associated with the left adrenal gland, along with a lack of division of the right adrenal gland. The segmentation maps obtained from TransAttUNet exhibit their limited capacity to learn complex representations corresponding to multiple organs, with varying shapes and sizes.
@@
Figure 4: Comparative performance of U-VixLSTM and other baseline architectures, on the ISIC dataset, through sample segmentation maps. The (a) input dermoscopic image, (b) corresponding ground truth, with the respective output from (c) U-VixLSTM, (d) Swin UNETR, (e) UNETR, (f) U-Net 3+, (g) TransAttUNet, and (h) DS-TransUNet.
@@
Table 3: Comparison with state-of-the-art models on the ISIC dataset. The best results are highlighted in bold.
@@
Swin UMamba 0.8237 0.7330 56.41
U-VixLSTM 0.85 0.7611 11.31
Table 4: Comparison with state-of-the-art models on the ACDC dataset. The best results are highlighted in bold.
@@
Model LVentricle RVentricle Myocardium mIOU mHD
UNet 0.7869 0.7956 0.891 0.713 7.29
V-Net 0.7273 0.6969 0.8108 0.6145 6.34
UNet++ 0.6155 0.6828 0.8302 0.5693 54.35
Attention U-Net 0.7087 0.7637 0.8689 0.6568 8.32
UNet with EfficientNet-b0 backbone 0.6236 0.5314 0.7069 0.5219 104.65
UNet with EfficientNet-b1 backbone 0.6543 0.5217 0.7299 0.5026 98.27
UNet 3+ 0.6903 0.6412 0.7355 0.5463 149.64
SwinUNet 0.8059 0.7741 0.8838 0.7076 6.02
UNETR 0.7052 0.6845 0.8316 0.6049 9.84
TransUNET 0.6398 0.7692 0.8714 0.6581 6.93
TransAttUNET 0.6338 0.742 0.8819 0.6432 95.35
Swin UMamba 0.664 0.7907 0.8699 0.672 5.45
DS-TransUNET 0.6396 0.7688 0.8654 0.654 5.32
UNETR++ 0.7026 0.6977 0.8264 0.654 5.32
U-VixLSTM 0.868 0.8345 0.9104 0.777 5.07
Figure 5: Comparative performance of U-VixLSTM and other baseline architectures, on the ACDC dataset, through sample segmentation maps. The (a) input CT image, (b) corresponding ground truth, with the respective output from (c) U-VixLSTM, (d) Swin UNETR, (e) UNETR, (f) V-Net, and (g) TransAttUNet
@@
Figure 7: Comparison between state-of-the-art models w.r.t number of parameters (in millions), TFLOPs, and memory size (in MB). Bubble size is indicative of model size.
Fig. 4 represents the sample segmentation output, as predicted by U-VixLSTM and other baseline models on the ISIC dataset. The prediction made by our U-VixLSTM exhibits higher accuracy and precision in relation to the ground truth masks, compared to the maps generated by the other baseline models. The sample outputs from the other state-of-the-art architectures mostly contain undersegmented or oversegmented regions. For example, along rows 2 and 3, the baseline models exhibit a significant amount of $FN$ and $FP$ regions; our U-VixLSTM, on the other hand, has a visually higher proportion of $TP$ regions and fewer $FP$ and $FN$ pixels. This indicates that these other models had difficulty accurately identifying the boundaries of the target region(s). The quantitative results, presented in Table 3, demonstrate (for our U-VixLSTM) the highest values of $DSC$ and $IoU$ along with the lowest $HD95$ metric score. This corroborates the qualitative observation of the enhanced segmentation accuracy of our proposed model.
Table 4 presents a comprehensive analysis of performance in different organs in the dataset ACDC concerning the DSC value. The average results for HD95 and IoU are tabulated for various cardiac organs. U-VixLSTM has outperformed baselines, attaining a DSC value of 86. 8%, 83. 45% and 91. 04% for the three distinct cardiac organs, respectively. The mean IoU is 77. 7%, exceeding the best baseline by approximately 6%. Fig. 5 shows sample segmentation maps generated by U-VixLSTM along with various baseline methods. The visual representation illustrates the higher-quality outputs generated by our proposed methods relative to established baselines. Instances of mis-segmentations and false positive pixels are evident in the sample maps from Swin UNETR, UNETR, and $V$-Net, as illustrated in Fig. 5(d)-(g).
@@
Fig. 6 illustrates sample feature maps obtained from the convolution and Vision-xLSTM modules within the feature extraction pathway. The maps visualize the representations acquired by the two modules of the proposed network in relation to the target structures, as specified in the ground truth. The shades of red represent the regions of strongest activation where the model allocates the majority of its attention, whereas the shades of dark blue signify the lowest activation levels. This visualization helps us observe the areas in the input image that are the most important to the model to define the target structure.
@@
The enhanced performance of U-VixLSTM in various modalities is due to its effective capture of local and contextual information. The CNN blocks initially hierarchically extract fine-grained details, such as edges and local patterns. The Vision-xLSTM blocks at the bottleneck connect distant sections of intermediate feature maps, effectively capturing the overall relationships and dependencies among various parts of target structures that have different shapes and sizes. This effectively constructs global contextual information. Incorporating skip connections that merge feature maps from the extraction path with the corresponding levels of the reconstruction path facilitates the localization of anatomical structures. Consequently, Vision-xLSTMs enhance the feature extraction process by acquiring superior feature representations while maintaining lower computational costs relative to baseline methods.

### v3

@@ 2.2 Feature reconstruction path
Input: Training dataset $\Delta=\{(I_{j},O_{j})\}_{j=1}^{m}$, Patch size $P$, Embedding dimension $Z$, Number of ViL blocks $L$, Projection matrix $K$, Positional embedding $K_{pos}$, learning rate $\epsilon$ and number of iterations $\iota$
@@
20:     for $i=1$ to $L$ do /* Apply ViL block */
21:         Compute memory cell state $\mathbf{C_{t}}$ and gate activations with normalizer state
22:         Update hidden state $h_{t}$
@@ 3 Experimental Details
This section describes the loss functions and performance metrics used for the evaluation, details of the datasets used and the experimental results. The PyTorch and MONAI frameworks were used to implement and train U-VixLSTM in Python 3.9 on an NVIDIA RTX A6000 GPU with 48GB of RAM. AdamW Optimizer was used, with a learning rate of 1e-4 and a weight decay of 1e-5.
@@ 3.1 Loss function and metrics
A combination of Dice $\mathcal{L}_{dice}=\gamma-\sum_{\gamma=1}^{c}\left(\frac{2\sum_{i=1}^{T}\hat{z}_{\gamma,i}z_{\gamma,i}+\mu}{\sum_{i=1}^{T}\hat{z}_{\gamma,i}+z_{\gamma,i}+\mu}\right)$ [5] and Categorical Cross Entropy [40] $\mathcal{L}_{cce}=-\frac{1}{T}\sum_{i=1}^{T}\sum_{\gamma=1}^{c}z_{\gamma,i}\log{(\hat{z}_{\gamma,i})}$ was employed to train the model. Here, $c$ represents the total number of classes to be predicted, $\hat{z}_{\gamma,i}$ and $z_{\gamma,i}$ correspond to the predicted and ground truth values (respectively) for the $i^{th}$ voxel for class $\gamma$. $T$ denotes the total number of voxels in the input and $\mu$ is the additive smoothing parameter used to overcome division-by-zero errors. The composite loss function $\mathcal{L}$ is expressed as
@@
Figure 3: Distribution of (a) various abdominal organs, (b) cardiac anatomical structures, and (c) skin lesion surface area, among different patients in the Synapse, ACDC and ISIC datasets, respectively.
@@ 3.2 Datasets
The data set Synapse comprises 30 CT volumes with sizes ranging from 512 $\times$ 512 $\times$ 85 to 512 $\times$ 512 $\times$ 198. CT volumes are manually annotated by experts in the field to highlight the different abdominal organs. A model is trained to segment nine distinct organs of the abdominal cavity, viz. spleen, left kidney, right kidney, liver, gall bladder, pancreas, stomach, right adrenal gland and left adrenal gland. The spleen, liver, and stomach are classified as larger organs, while the kidneys, gall bladder, pancreas, and adrenal glands are smaller in size. The CT volumes were divided into training and test sets, with 80% used for training and the remaining 20% used for testing to evaluate the generalization performance of the model. The voxel intensities were restricted between -170 HU and 250 HU, followed by normalization of the intensity to the range of [0,1]. Data augmentation was used to expand the size of the training dataset.
@@
Fig. 3 illustrates the distribution of abdominal and cardiac organ volumes, as well as the surface area of skin lesions, between patients in the three datasets. The plots demonstrate substantial evidence of inter-patient variability across the datasets, indicating significant differences in the sizes of various organs or lesion areas among patients. Quite a substantial difference in morphology is evident in the spleen, liver, and stomach, as shown in Fig. 3(a), in the cardiac organs of Fig. 3(b), and in the total surface area of the lesion illustrated in Fig. 3(c), as attributed to long tails and broader distributions. The median values indicate significant heterogeneity in the average volumes of the various abdominal organs. The liver exhibits the largest average volume, followed by the stomach and spleen. The adrenal gland exhibits the lowest average volume among the glands. The diverse sizes of target structures hinder the ability of the model to generalize across datasets. Fig. 3(c) illustrates a right-skewed distribution of the surface area of the lesions. This suggests that patients generally have smaller lesion areas compared to larger ones. This may indicate a class imbalance within the dataset.
@@
3.3 Results and Discussion
Table 1: Comparison of different variants of U-VixLSTM with increasing number of ViL blocks (×$\times L$) and convolution layers, on Synapse data. Best results and the selected model configuration are highlighted in bold.
@@ 3.3 Results and Discussion
Table 1 presents an ablation study on Synapse data, demonstrating the impact of different numbers ($L$) of ViL blocks and a varying number of convolution layers along the feature extraction path. The average scores of $DSC$, $HD95$, and $IoU$ are reported for the different organs. The ablation study reflects the trade-off between the complexity of the model and its performance. The best $DSC$ and $IoU$ values were observed for 6 ViL blocks. The performance degrades when increasing the number of ViL blocks beyond 6, which indicates overfitting. The model appears to memorize the training data instead of learning generalizable features, which occurs as the number of parameters increases due to the addition of more ViL blocks. The 6-block configuration appears to be sufficient on this dataset to capture the global context without the risk of overfitting. Similarly, increasing the number of convolution layers from 3 to 4 shows a significant gain in metric scores. This indicates the importance of hierarchical depth in extracting rich representations. However, the performance saturates with a further increase in the number of layers, which indicates minimal benefits with an increase in the computational load. The two-part ablation study empirically justifies the design choice of U-VixLSTM to extract robust spatial features hierarchically, followed by modeling global context. This approach efficiently mitigates the risk of overfitting while ensuring computational feasibility.
Table 2: Comparison with state-of-the-art models on multi-organ segmentation (Synapse) dataset, with best results marked in bold.
@@
$U$-Net 0.9112 0.9007 0.9181 0.5645 0.9572 0.6967 0.7800 0.6160 0.5872 0.7702 0.6667 38.83
$V$-Net 0.8874 0.9251 0.9244 0.5858 0.9487 0.7511 0.7953 0.6004 0.4497 0.7631 0.6537 25.29
$U$-Net++ 0.9118 0.9196 0.8905 0.6921 0.9524 0.7536 0.7869 0.5245 0.1032 0.7261 0.6254 55.27
Attention $U$-Net 0.9109 0.8770 0.8720 0.5835 0.9585 0.5566 0.7846 0.5991 0.5398 0.7424 0.6262 51.61
$U$-Net + EfficientNet-b0 0.8541 0.8919 0.8804 0.6370 0.9077 0.5018 0.7097 0.6048 0.5195 0.7230 0.6056 67.47
$U$-Net + EfficientNet-b1 0.7414 0.8059 0.8256 0.5500 0.9019 0.5416 0.6160 0.5250 0.5035 0.6679 0.5539 75.85
$U$-Net 3+ 0.8736 0.9229 0.8584 0.6659 0.9383 0.6823 0.7259 0.5657 0.4764 0.7455 0.6254 54.43
Swin UNETR 0.9482 0.9300 0.9245 0.7617 0.9622 0.8046 0.8059 0.6890 0.6159 0.8269 0.7261 13.99
UNETR 0.8951 0.9055 0.8950 0.5274 0.9461 0.6668 0.7847 0.5268 0.4978 0.7384 0.6284 26.94
TransUNet 0.8520 0.8828 0.7850 0.7608 0.5870 0.7218 0.8773 0.5876 0.4672 0.7246 0.6679 36.81
TransAttUNet 0.9045 0.8761 0.9160 0.4958 0.9408 0.5997 0.6682 0.6559 0.5948 0.7391 0.6580 15.82
Swin-UMamba 0.8778 0.9037 0.9112 0.4206 0.9161 0.6238 0.6527 0.5890 0.5617 0.7174 0.6334 20.10
UNETR++ 0.8061 0.8050 0.8201 0.5208 0.8868 0.5249 0.5576 0.5890 0.4577 0.6631 0.5340 8.61
DS-TransUNet 0.4965 0.6113 0.5252 0.5397 0.6790 0.5295 0.5857 0.5745 0.5266 0.5631 0.4644 14.79
U-VixLSTM 0.9500 0.9371 0.9366 0.8104 0.9635 0.7878 0.8304 0.6709 0.6458 0.8318 0.7286 4.80
@@
The performance of our U-VixLSTM was next compared with that of the state-of-the-art (SOTA) algorithms in the literature in terms of metrics $DSC$, $IoU$, $HD95$ on Synapse, ISIC and ACDC datasets. Table 2 presents a comprehensive breakdown of performance for different organs (in Synapse data) in the context of DSC. The average results for $HD95$ and IoU are provided for the nine abdominal organs. U-VixLSTM is found to outperform other SOTA, with respect to the mean DSC, IoU and $HD95$, with scores of 83.18%, 72.86%, and 4.8, respectively. Our proposed U-VixLSTM demonstrated superior performance in generating the highest $DSC$ values for the segmentation of larger organs (such as the spleen, liver) and smaller organs (such as the kidneys, pancreas, gall bladder and left adrenal gland). Performance remained stable, despite the reduction in organ size, as indicated by the consistently high $DSC$ scores observed in smaller and larger organs. This illustrates the robustness of our model in utilizing acquired knowledge on anatomical structures with varying shapes and sizes. It is found to precisely identify and define the target areas possessing unique structures.
Figure 4: Dot plot of (a) DSC, (b) IoU, and (c) $HD95$ metrics for evaluating the performance of U-VixLSTM against other baselines on Synapse dataset. The radius of the circles indicates the metric values, while the color intensity signifies the standard deviation.
@@
Figure 5: Comparative performance of U-VixLSTM with other baseline architectures, on the Synapse dataset, through sample segmentation maps. The first row in each block represents a sample CT slice. The second row in each block provides zoomed-in boxes to provide a magnified view of specific regions. (a) Input CT image, (b) corresponding ground truth, with the respective output from (c) U-VixLSTM, (d) Swin UNETR, (e) UNETR, (f) V-Net, and (g) TransAttUNet.
Fig. 4 visualizes performance consistency across different metrics for U-VixLSTM and baselines concerning different abdominal organs. The visualization represents the metric values (denoted by the dot radius) and consistency (represented by the color intensity of the dots). U-VixLSTM exhibits consistent superiority across various organs, as evidenced by the larger radii and uniformly pale color intensities of the dots in Fig. 4(a) and (b). This signifies higher accuracy with low standard deviation. Therefore, the high accuracy of the proposed model is both consistent and reliable. Conversely, Swin-UMamba, DS-TransUNet, TransAttUNet and TransUNet, having darker and smaller dots, signifies high variability and lesser accuracy than U-VixLSTM. Fig. 4(c) represents the $HD95$ metric values across all the organs for different models. A smaller dot radius means higher accuracy in delineating the boundaries for this metric. U-VixLSTM consistently demonstrates smallest dot radii with pale color intensities for nearly all the organs. This indicates the superior boundary delineating capabilities compared to baselines, like U-Net 3+, U-Net with EfficientNet backbone, Attention U-Net and TransUNet, having significantly larger and darker dots. While TransUNet attains the highest mean DSC value for the stomach, it exhibits a considerably greater standard deviation and higher $HD95$ values in comparison to U-VixLSTM. This indicates that while TransUNet effectively segments the majority of the stomach region, it encounters difficulties in accurately delineating the organ boundary and sustaining consistent performance.
@@
Fig. 5 shows the sample segmentation output of U-VixLSTM and other architectures, such as Swin UNETR, UNETR, $V$-Net, and TransAttUNet, on the Synapse dataset. It is evident that the U-VixLSTM model demonstrates a significantly higher level of similarity to the ground truth, compared to the maps produced by these other baseline models. The results of the Swin UNETR and UNETR models show $FP$ regions corresponding to the stomach, in the second row of Figs. 5(d)-(e). The results of $V$-Net demonstrate a lack of precision in identifying the anatomical regions associated with the adrenal gland, kidneys, pancreas and stomach. The segmentation maps obtained from TransAttUNet exhibit their limited capacity to learn complex representations corresponding to multiple organs, with varying shapes and sizes.
The hybrid architecture of U-VixLSTM shows effectiveness in segmenting abdominal organs with diverse shapes and sizes. The CNN layers effectively capture intricate details in high-resolution maps, which is essential for accurately delineating small-sized organs. The ViL block strongly models the global context, making it well-suited for the segmentation of larger organs.
@@
Figure 6: Comparative performance of U-VixLSTM and other baseline architectures, on the ISIC dataset, through sample segmentation maps. (a) Input dermoscopic image, (b) corresponding ground truth, with the respective output from (c) U-VixLSTM, (d) Swin UNETR, (e) UNETR, (f) U-Net 3+, (g) TransAttUNet, and (h) DS-TransUNet.
@@
Table 3: Comparison with state-of-the-art models on the ISIC dataset, with best results marked in bold.
@@
Swin-UMamba 0.8237 0.7330 56.41
[…]
U-VixLSTM 0.8500 0.7611 11.31
Fig. 6 represents the sample segmentation output, as predicted by U-VixLSTM and other baseline models on the ISIC dataset. The prediction made by our U-VixLSTM exhibits higher accuracy and precision in relation to the ground truth masks, compared to the maps generated by the other baseline models. The sample outputs from the other state-of-the-art architectures mostly contain undersegmented or oversegmented regions. For example, along rows 2 and 3, the baseline models exhibit a significant amount of $FN$ and $FP$ regions; our U-VixLSTM, on the other hand, has a visually higher proportion of $TP$ regions and fewer $FP$ and $FN$ pixels. This indicates that these other models had difficulty accurately identifying the boundaries of the target region(s). The quantitative results, presented in Table 3, demonstrate the highest values of $DSC$ and $IoU$ along with the lowest $HD95$ metric score for our U-VixLSTM. This corroborates the qualitative observation of the enhanced segmentation accuracy of our proposed model.
Figure 7: Dot plot of (a) DSC, (b) IoU, and (c) $HD95$ metrics for evaluating the performance of U-VixLSTM against other baselines on ISIC data. The radius of the circles indicates the metric values, while the color intensity signifies the standard deviation.
Fig. 7 provides a detailed analysis of U-VixLSTM against the baselines across all the metrics. The pale blue hue of the dot corresponding to U-VixLSTM across all the subplots signifies a minimal standard deviation. This demonstrates the superior performance of the model across the diverse dermoscopic images. Precisely delineating the boundaries of the affected region is a significant challenge in skin lesion analysis. Fig. 7(c) demonstrates the efficacy of U-VixLSTM in addressing this challenge, as indicated by the smallest dot representing the minimum $HD95$ value of 11.31. On the contrary, Transformer or Mamba-based models exhibit a larger radius (suboptimal boundary delineation), with darker blue hues signifying discrepancies in boundary prediction.
Table 4: Comparison with state-of-the-art models on the ACDC dataset, with best results marked in bold.
@@
Model LVentricle RVentricle Myocardium Mean mIOU mHD95
$U$-Net 0.7869 0.7956 0.8910 0.8245 0.7130 7.29
$V$-Net 0.7273 0.6969 0.8108 0.7450 0.6145 6.34
$U$-Net++ 0.6155 0.6828 0.8302 0.7095 0.5693 54.35
Attention $U$-Net 0.7087 0.7637 0.8689 0.7804 0.6568 8.32
$U$-Net + EfficientNet-b0 backbone 0.6236 0.5314 0.7069 0.6206 0.5219 104.65
$U$-Net + EfficientNet-b1 backbone 0.6543 0.5217 0.7299 0.6353 0.5026 98.27
$U$-Net 3+ 0.6903 0.6412 0.7355 0.6890 0.5463 149.64
Swin UNETR 0.8059 0.7741 0.8838 0.8213 0.7076 6.02
UNETR 0.7052 0.6845 0.8316 0.7404 0.6049 9.84
TransUNet 0.6398 0.7692 0.8714 0.7601 0.6581 6.93
TransAttUNet 0.6338 0.7420 0.8819 0.7526 0.6432 95.35
Swin-UMamba 0.6640 0.7907 0.8699 0.7749 0.6720 5.45
DS-TransUNet 0.6396 0.7688 0.8654 0.7579 0.6540 5.32
UNETR++ 0.7026 0.6977 0.8264 0.7422 0.6121 6.59
U-VixLSTM 0.8680 0.8345 0.9104 0.8710 0.7770 5.07
Figure 8: Dot plot of (a) DSC, (b) IoU, and (c) $HD95$ metrics for evaluating the performance of UVixLSTM against other baselines on ACDC data. The radius of the circles indicates the metric values, while the color intensity signifies the standard deviation.
Figure 9: Comparative performance of U-VixLSTM and other baseline architectures, on the ACDC dataset, through sample segmentation maps. The first row in each block represents a sample CT slice. The second row in each block provides zoomed-in boxes to provide a magnified view of specific regions. The (a) input CT image, (b) corresponding ground truth, with the respective output from (c) U-VixLSTM, (d) Swin UNETR, (e) UNETR, (f) V-Net, and (g) TransAttUNet
Table 4 presents a comprehensive analysis of performance in different organs in the dataset ACDC concerning the DSC value. The average results for HD95 and IoU are tabulated for various cardiac organs. U-VixLSTM has outperformed baselines, attaining a DSC value of 86.8%, 83.45% and 91.04% for the three distinct cardiac organs, respectively. The mean IoU is 77.7%, exceeding the best baseline by approximately 6%. The dotplot in Fig. 8 provides a detailed analysis of the quantitative results. U-VixLSTM consistently displays dots with larger radii and pale color intensities across all the cardiac components in Fig. 8(a) and (b). This illustrates a balanced and high accuracy across the multiple interconnected cardiac structures. Other approaches, such as DS-TransUNet, Swin U-Mamba, TransUNet and Swin UNETR demonstrate either comparable or lower $HD95$ value, as evidenced from Fig. 8(c). However, their performance in terms of DSC and IoU is either significantly lower, as shown in Table 4 and Fig. 8(a) and (b) or have a higher standard deviation concerning $HD95$ (Fig. 8(c)) compared to U-VixLSTM. This suggests U-VixLSTM achieves improved and consistent performance in both overlap-based and boundary-based metrics. The ViL block at the bottleneck of U-VixLSTM efficiently captures the global view of the cardiac anatomy. This intermediate output guides the feature reconstruction path to generate output that is accurate regarding boundary delineation and region overlap.
Fig. 9 shows sample segmentation maps generated by U-VixLSTM along with various baseline methods. The visual representation illustrates the higher-quality outputs generated by our proposed methods relative to established baselines. Instances of mis-segmentations and false positive pixels are evident in the sample maps from Swin UNETR, UNETR, and $V$-Net, as illustrated in Fig. 9(d)-(g).
@@
Fig. 10 illustrates sample feature maps obtained from the convolution and ViL modules within the feature extraction pathway. The maps visualize the representations acquired by the two modules of the proposed network in relation to the target structures, as specified in the ground truth. The shades of red represent the regions of strongest activation where the model allocates the majority of its attention, whereas the shades of dark blue signify the lowest activation levels. This visualization helps us observe the areas in the input image that are the most important to the model to define the target structure.
@@
The enhanced performance of U-VixLSTM in various modalities is due to its effective capture of local and contextual information. The CNN blocks initially hierarchically extract fine-grained details, such as edges and local patterns. The ViL blocks in the bottleneck connect distant sections of intermediate feature maps, effectively capturing the general relationships and dependencies between various parts of target structures that have different shapes and sizes. This effectively constructs global contextual information. Incorporating skip connections that merge feature maps from the extraction path with the corresponding levels of the reconstruction path facilitates the localization of anatomical structures. Consequently, ViL enhance the feature extraction process by acquiring superior feature representations while maintaining lower computational costs relative to baseline methods. The gating mechanism of ViL selectively updates the memory matrix to store sharp transitions corresponding to the organ boundaries. This explains the superiority of U-VixLSTM in efficiently modelling the boundaries of diverse anatomical structures in comparison to Transformer-based models. Transformer computes the contextual embedding of each patch using a weighted average of all the image patches. Although being able to effectively capture the global context, the averaging step smoothens the fine-grained details necessary for capturing the boundaries.
Figure 11: Comparison between state-of-the-art models with respect to number of parameters (in millions), TFLOPs, and model size on disk (in MB). Bubble size is indicative of model size.